OpenStudy (anonymous):

If the area of a rectangle 24a2b and the length is 8ab2, what would the width of the rectangle, given that width is found by dividing area by length? Simplify the answer. the quantity of 3 times a all over b b over the quantity 3 times a 3 over the quantity a times b 2st

OpenStudy (anonymous):

PLEASE HELP

OpenStudy (mokeira):

 \[Area=Length \times Width\] \[Area=24a ^{2}b\] \[Length=8ab ^{2}\] \[24a ^{2}b=8ab ^{2} \times width\] \[\frac{ 24a ^{2}b }{8ab ^{2} }=width\] simplify \[\frac{ 3a }{ b }=width \] \[3ab ^{-1}= width\]

OpenStudy (mokeira):

This matches the first choice

OpenStudy (anonymous):

whats the answer

OpenStudy (anonymous):

thanks

OpenStudy (mokeira):

U welcome :)